OpenStudy (anonymous):

-10+√x+8= -4 A. -2 B. 36 C. 44 D. 28 Please help. I'm absolutely terrible at Algebra 2 :(

OpenStudy (anonymous):

It's asking for the solution of the equation.

OpenStudy (anonymous):

Clarification: does your equation actually look like this? \[-10 + \sqrt{x +8} = -4\]

OpenStudy (usukidoll):

No wonder I was getting weird results when I did this in my head. Whatever you wrote @Cardinal_Carlo makes better sense

OpenStudy (anonymous):

@UsukiDoll I felt your struggles also

OpenStudy (usukidoll):

mhm the formatting of the equation in question makes a very big difference.

OpenStudy (anonymous):

I'm sorry, guys. I'm new to this website. I never know how to type it out online. Sorry it's so frustrating :( But yes, that's what it is @Cardinal_Carlo @UsukiDoll

OpenStudy (anonymous):

That's okay @babyxfce. You're doing just fine. We're just asking for some clarity. we first group our constants: \[\sqrt{x + 8} = -4 + 10\] then we square both sides of the equation like so: \[x + 8 = 6^{2}\] we simplifiy and get: \[x = 36 - 8 = 28\] which happens to be D